Olsen Outfitters Inc. believes that its optimal capital structure consists of 60% common equity and 40% debt, and its tax rate is 40%. Olsen must raise additional capital to fund its upcoming expansion. The firm will have $2 million of retained earnings with a cost of rs = 14%. New common stock in an amount up to $10 million would have a cost of re = 16%. Furthermore, Olsen can raise up to $4 million of debt at an interest rate of rd = 9% and an additional $5 million of debt at rd = 12%. The CFO estimates that a proposed expansion would require an investment of $5.8 million. What is the WACC for the last dollar raised to complete the expansion? Round your answer to two decimal places.
Answer #1

equity = 60% of amount to raise = 0.6*5.8 = 3.48bn

debt = 40% of capital to raise =0.4*5.8 = 2.32 bn

company will have to raise equity as retained earnings of 2m is not sufficient

company can raise debt in 1st tranche as 4 ml limit for debt tranche 1 is sufficient

After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 9*(1-0.4)
= 5.4
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=5.4*0.4+16*0.6
WACC =11.76%